# Detect Anything 3D in the Wild

### v2

@@ 3.4 Loss
Our loss function comprises three components, the depth loss $\mathcal{L}_{\text{depth}}$, the camera intrinsic loss $\mathcal{L}_{\text{cam}}$, and the detection loss $\mathcal{L}_{\text{det}}$. The overall loss is defined as the sum of these three components. For depth loss $\mathcal{L}_{\text{depth}}$, we adopt the commonly used SILog loss [19, 64] to supervise depth prediction. For camera intrinsic loss $\mathcal{L}_{\text{cam}}$, we follow the dense camera ray approach [54, 30] to represent intrinsics and also employ the SILog loss to measure deviations between predicted and ground-truth parameters. At last, for detection loss $\mathcal{L}_{\text{det}}$, we use the smooth L1 loss [80, 40, 66] to regress 3D bounding boxes parameters and predicted IOU scores and the Chamfer loss [8, 74] for rotation matrices. Detailed formulations of these loss functions can be found in Section 7.3.
@@ 3.5 Prompt Interaction
Text Prompts. Recent 2D foundation models like Grounding DINO [44] are able to detect bounding box for the open-vocabulary object specified by users using text prompt. DetAny3D can further generate 3D bouding box using the prediction of Grounding DINO, which enables text as prompts in the zero-shot interface.
@@ 4 Experiment
- Cube R-CNN [8] - - - 8.22 9.43 -
Cube R-CNN OVMono3D [74] - - - 4.97 10.89 -
@@
$\Delta$ +5.36 +4.28 -
Grounding DINO OVMono3D [74] 4.71 / 4.71 4.07 / 16.78 13.21 / 13.21 5.88 / 10.98 9.20 / 10.27 0.37 / 8.48
@@
Ground Truth OVMono3D [74] 8.44 17.16 14.12 10.06 10.23 18.05
@@ 4.1 Experimental Setup
DA3D Benchmark. We present DA3D, a unified 3D detection dataset that aggregates 16 diverse datasets for 3D detection and depth estimation. Building upon Omni3D’s original datasets (Hypersim [57], ARKitScenes [4], Objectron [2], SUNRGBD [61], KITTI [24], and nuScenes [9]), we incorporate additional four outdoor detection datasets (Argoverse2 [70], A2D2 [25], Waymo [62], Cityscapes3D [21]), one indoor detection dataset (3RScan [65]), and five depth and intrinsic datasets (Scannet [17], Taskonomy [77], DrivingStereo [71], Middlebury [59], IBIMS-1 [34]). All data is standardized with monocular images, camera intrinsics, 3D bounding boxes, and depth maps. Following prior work [74], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as zero-shot test classes. We select Cityscapes3D, Waymo, and 3RScan as our zero-shot datasets with novel camera configurations, where 3RScan also contains novel object categories. Depth supervision from LiDAR, RGB-D, and stereo sensors enhances 75% of training samples, while intrinsic parameters cover 20 camera configurations across 0.4 million frames (2.5× Omni3D’s scale). Dataset statistics and splits are detailed in Section 6. All data are subject to their respective licenses.
@@
Baselines. We choose Cube R-CNN [8] and OVMono3D [74] as our primary baselines, as their settings align most closely with our experimental protocol: Cube R-CNN is a benchmark provided by the Omni3D dataset. It is a unified detector capable of performing detection on predefined categories. OVMono3D is a recently available open-vocabulary 3D detector on the Omni3D dataset. It lifts 2D detection to 3D by connecting the open-vocabulary 2D detector Grounding DINO [44] with a detection head.
@@
Metrics. We adopt the metrics in the Omni3D benchmark [8], which is Average Precision (AP). Predictions are matched to ground-truth by measuring their overlap using IoU3D, which computes the intersection-over-union (IoU) of 3D cuboids. The IoU3D thresholds range from $\tau\in[0.05,0.10,\dots,0.50]$. For experiments using text prompts, we additionally employ target-aware metrics from OVMono3D [74]: Prompt the detector only with category names present in the per-image annotations instead of providing an exhaustive category list. This addresses severe naming ambiguity (e.g., ”trash can” vs. ”rubbish bin”) and missing annotation issues prevalent in indoor datasets like 3RScan (see Section 8.).
@@
Implementation Details. We implement DetAny3D via PyTorch [53]. We use the pretrained ViT-L DINOv2 [51, 54] and ViT-H SAM [33] as our initial models, with SAM serving as the promptable backbone, where the encoder is frozen during training. All main experiments are conducted using 8 NVIDIA A100 machines with 8 GPUs for each and a batch size of 64. The model is trained for 80 epochs, taking approximately 2 weeks to complete. The training uses the AdamW [47] optimizer with an initial learning rate of 0.0001, adjusted according to the cosine annealing policy [46]. During box prompt training, we apply a 0.1 positional offset disturbance. For point prompt training, points are randomly selected from the mask. Text prompts are converted into box prompts via Grounding DINO SwinT [44]. For fair comparisons, all baseline-related experiments incorporate intrinsic prompts and use aligned prompt inputs.
@@
ImVoxelNet [58] 23.5 23.4 21.5 - - - - - - 9.4
SMOKE [45] 25.9 20.4 20.0 - - - - - - 10.4
OV-Uni3DETR [68] 35.1 33.0 31.6 - - - - - - -
Cube R-CNN [8] 36.0 32.7 31.9 32.50 30.06 15.33 41.73 50.84 7.48 23.26
@@
[…]
@@ 4.2 Main Results
Zero-shot Category Performance. In this experiment, we use two sources for the prompt input: text prompt processed by Grounding DINO and box prompt from ground-truth 2D bounding box. We evaluate our model on KITTI, SUNRGBD, and ARKitScenes datasets with the same zero-shot categories as OVMono3D [74]. As shown in Table 1 (left), our DetAny3D demonstrates superior zero-shot adaptation performance compared to the OVMono3D baseline. When using Grounding DINO for text prompt input, our method achieves significant improvements of 21.02 ${\rm AP_{3D}}$ on KITTI, 4.29 ${\rm AP_{3D}}$ on SUNRGBD, and 11.35 ${\rm AP_{3D}}$ on ARKitScenes under the target-aware metric. When using 2D ground-truth as box prompt input, DetAny3D attains 28.96 ${\rm AP_{3D}}$ on KITTI, 39.09 ${\rm AP_{3D}}$ on SUNRGBD, and 57.72 ${\rm AP_{3D}}$ on ARKitScenes, showing 3.4×, 2.3×, and 4.1× gains over the baseline, respectively. This substantial performance gap highlights our method’s enhanced ability to generalize to novel object categories.
Zero-shot Camera Performance. To assess robustness against novel camera parameters, we conduct cross-dataset evaluation as shown in Table 1 (right). For Cityscapes3D and Waymo, We use Cube R-CNN’s 2D detections and ground-truth as box prompt and Grounding DINO processed text prompt for comparison. For 3RScan, due to namespace inconsistency with Cube R-CNN’s predefined categories and the presence of novel classes, we only use text prompt and ground-truth box prompts, benchmarking against OVMono3D. DetAny3D exhibits strong adaptation to unseen camera configurations. When using Cube R-CNN-aligned prompts, our model achieves $\rm{AP_{3D}}$ scores of 10.33 and 15.17 on Cityscapes3D and Waymo, respectively, surpassing Cube R-CNN by +2.11 and +5.74. With text prompts, under identical settings as OVMono3D [74], our method improves $\rm{AP_{3D}}$ by +4.73 on Cityscapes3D, +5.68 on Waymo, and +1.1 on 3RScan under target-aware metrics. Both models show low scores on conventional metrics for 3RScan due to severe naming ambiguity and missing annotations. Using 2D ground-truth as box prompts, DetAny3D attains $\rm{AP_{3D}}$ of 16.88, 15.83, and 21.36 across the three datasets, outperforming OVMono3D by +6.82, +5.6, and +3.31, respectively. These results highlight the effectiveness of our architecture and its potential for real-world applications with arbitrary camera configurations.
@@
Figure 4: Qualitative Results. We present qualitative examples from open-world detection. In each pair of images, the top row is produced by OVMono3D, and the bottom row by DetAny3D. For each example, the left sub-figure overlays the projected 3D bounding boxes, while the right sub-figure shows the corresponding bird’s-eye view with 1m×$\times$1m grids as the background.
In-domain Performance We also evaluate our model’s in-domain detection capability using two prompt sources: 2D detections from Cube R-CNN and 2D ground-truth. In addition to the unified model, we also train our model on Omni3D_out for comparison. As shown in Table 2, DetAny3D achieves competitive detection results with Cube R-CNN when provided with aligned input. Moreover, when using GT as 2D prompts, DetAny3D significantly outperforms OVMono3D, with an overall $\rm{AP_{3D}}$ improvement of 9.06 on Omni3D. This performance gap suggests that when Cube R-CNN is used as the 2D input, the limitations of Cube R-CNN partially constrain the performance of our model. By matching with stronger 2D prompts, our model has the potential for even higher performance.
@@ 4.3 Possible Applications of DetAny3D
3D Bounding Box Guided Video Generation. We feed the outputs of DetAny3D into Sora to achieve zero-shot 3D bounding box guided video generation in open-world settings. As illustrated in Figure 3, we compare three prompting strategies: (i) image + 3D box + text control, (ii) image + 2D box + text control, and (iii) image + text control. With 3D bounding box constraints, Sora produces videos that more closely align with the intended descriptions.
@@ 4.4 Ablation Studies
Table 3: Ablation study of DetAny3D. The table shows the impact of different design choices on the ${\rm AP_{3D}}$ performance. Each component is progressively added to the model. To save computational resources, ablation studies are conducted on 10% of the full training dataset.
@@
As shown in Table 3, we conduct ablation studies on key components of DetAny3D, illustrating the evolution from a vanilla SAM-based baseline to the well developed DetAny3D capable of extracting generalizable 3D features. The base model extends SAM by introducing 3D box tokens and a 3D prediction head, enabling direct 3D bounding box estimation. Additional ablations, including backbone choices and prompt types, are presented in Section 9.
@@
  • Effectiveness of ZEM. ZEM mechanism integrate geometric features through zero-initialized layers, which reduces catastrophic forgetting in model training across datasets with varying camera parameters, scenes, and depth distributions.
@@ 4.5 Qualitative Results
We provide qualitative results in comparison with OVMono3D. Both methods are driven by text prompts, while Grounding DINO is used as the 2D detector. As shown in Figure 4, our model can predict more accurate intrinsic when the camera intrinsics are unknown and infer consistent camera parameters and detection outputs.
@@ Acknowledgements
We sincerely thank Jiazhi Yang, Tianyu Li, Haochen Tian, Jisong Cai, and Li Chen for their invaluable discussions and constructive feedback throughout this project. Their insights and expertise have contributed significantly to the success of this work. We also appreciate the continuous support and encouragement from all the members of OpenDriveLab. This work is supported by the National Key Research and Development Program of China (2024YFE0210700) and NSFC (62206172). The project is in part financially supported by Meituan Inc.
@@ 6 DA3D
DA3D is a unified 3D detection dataset, consists of 16 diverse datasets. It builds upon six datasets in Omni3D—Hypersim [57], ARKitScenes [4], Objectron [2], SUNRGBD [61], KITTI [24], and nuScenes [9]—while partially incorporating an additional 10 datasets to further enhance the scale, diversity, and generalization capabilities of 3D detection models. As shown in Figure 5, DA3D comprises 0.4 million frames (2.5× the scale of Omni3D), spanning 20 distinct camera configurations.
The dataset is standardized with the similar structure to Omni3D [8], including monocular RGB images, camera intrinsics, 3D bounding boxes, and depth maps. DA3D is designed to test 3D detection models across a wide variety of environments, camera configurations, and object categories, offering a more comprehensive evaluation setting.
@@ 6.1 Dataset Composition
For training and evaluation, we follow the dataset splitting strategy used in prior works [8]. Specifically:
@@ 6.3 Evaluation Setup
Zero-Shot Categories. Following prior work [74], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as unseen classes for zero-shot testing.

### v3

@@ 3.4 Loss
Our loss function comprises three components, the depth loss $\mathcal{L}_{\text{depth}}$, the camera intrinsic loss $\mathcal{L}_{\text{cam}}$, and the detection loss $\mathcal{L}_{\text{det}}$. The overall loss is defined as the sum of these three components. For depth loss $\mathcal{L}_{\text{depth}}$, we adopt the commonly used SILog loss [NIPS2014_7bccfde7, Ummenhofer_2017_CVPR] to supervise depth prediction. For camera intrinsic loss $\mathcal{L}_{\text{cam}}$, we follow the dense camera ray approach [piccinelli2024unidepth, he2024diffcalib] to represent intrinsics and also employ the SILog loss to measure deviations between predicted and ground-truth parameters. At last, for detection loss $\mathcal{L}_{\text{det}}$, we use the smooth L1 loss [zhang2023monodetr, li2024unimode, wang2021fcos3d] to regress 3D bounding boxes parameters and predicted IOU scores and the Chamfer loss [brazil2023omni3d, yao2024open] for rotation matrices. Detailed formulations of these loss functions can be found in the supplementary material Section 7.3.
@@ 3.5 Prompt Interaction
Text Prompts. Recent 2D foundation models like Grounding DINO [liu2024grounding] are able to detect bounding box for the open-vocabulary object specified by users using text prompt. DetAny3D can further generate 3D bouding box using the prediction of Grounding DINO, which enables text as prompts in the zero-shot interface.
@@ 4 Experiment
- Cube R-CNN [brazil2023omni3d] - - - 8.22 9.43 -
Cube R-CNN OVMono3D [yao2024open] - - - 4.97 10.89 -
@@
$\Delta$ - - - +5.36 +4.28 -
Grounding DINO OVMono3D [yao2024open] 4.71 / 4.71 4.07 / 16.78 13.21 / 13.21 5.88 / 10.98 9.20 / 10.27 0.37 / 8.48
@@
Ground Truth OVMono3D [yao2024open] 8.44 17.16 14.12 10.06 10.23 18.05
@@ 4.1 Experimental Setup
DA3D Benchmark. We present DA3D, a unified 3D detection dataset that aggregates 16 diverse datasets for 3D detection and depth estimation. Building upon Omni3D’s original datasets (Hypersim [roberts2021hypersim], ARKitScenes [baruch1arkitscenes], Objectron [ahmadyan2021objectron], SUNRGBD [song2015sun], KITTI [geiger2013vision], and nuScenes [caesar2020nuscenes]), we incorporate additional four outdoor detection datasets (Argoverse2 [wilson2argoverse], A2D2 [geyer2020a2d2], Waymo [sun2020scalability], Cityscapes3D [gahlert2020cityscapes]), one indoor detection dataset (3RScan [wald2019rio]), and five depth and intrinsic datasets (Scannet [dai2017scannet], Taskonomy [zamir2018taskonomy], DrivingStereo [yang2019drivingstereo], Middlebury [scharstein2002taxonomy], IBIMS-1 [koch2018evaluation]). All data is standardized with monocular images, camera intrinsics, 3D bounding boxes, and depth maps. Following prior work [yao2024open], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as zero-shot test classes. We select Cityscapes3D, Waymo, and 3RScan as our zero-shot datasets with novel camera configurations, where 3RScan also contains novel object categories. Depth supervision from LiDAR, RGB-D, and stereo sensors enhances 75% of training samples, while intrinsic parameters cover 20 camera configurations across 0.4 million frames (2.5× Omni3D’s scale). Dataset statistics and splits are detailed in Supplementary material Section 6. All data are subject to their respective licenses.
@@
Baselines. We choose Cube R-CNN [brazil2023omni3d] and OVMono3D [yao2024open] as our primary baselines, as their settings align most closely with our experimental protocol: Cube R-CNN is a benchmark provided by the Omni3D dataset. It is a unified detector capable of performing detection on predefined categories. OVMono3D is a recently available open-vocabulary 3D detector on the Omni3D dataset. It lifts 2D detection to 3D by connecting the open-vocabulary 2D detector Grounding DINO [liu2024grounding] with a detection head.
@@
Metrics. We adopt the metrics in the Omni3D benchmark [brazil2023omni3d], which is Average Precision (AP). Predictions are matched to ground-truth by measuring their overlap using IoU3D, which computes the intersection-over-union (IoU) of 3D cuboids. The IoU3D thresholds range from $\tau\in[0.05,0.10,\dots,0.50]$. For experiments using text prompts, we additionally employ target-aware metrics from OVMono3D [yao2024open]: Prompt the detector only with category names present in the per-image annotations instead of providing an exhaustive category list. This addresses severe naming ambiguity (e.g., ”trash can” vs. ”rubbish bin”) and missing annotation issues prevalent in indoor datasets like 3RScan (see Supplementary material Section 8.).
@@
Implementation Details. We implement DetAny3D via PyTorch [paszke2019pytorch]. We use the pretrained ViT-L DINOv2 [oquab2024dinov2, piccinelli2024unidepth] and ViT-H SAM [kirillov2023segment] as our initial models, with SAM serving as the promptable backbone, where the encoder is frozen during training. All main experiments are conducted using 8 NVIDIA A100 machines with 8 GPUs for each and a batch size of 64. The model is trained for 80 epochs, taking approximately 2 weeks to complete. The training uses the AdamW [loshchilov2017decoupled] optimizer with an initial learning rate of 0.0001, adjusted according to the cosine annealing policy [loshchilov2016sgdr]. During box prompt training, we apply a 0.1 positional offset disturbance. For point prompt training, points are randomly selected from the mask. Text prompts are converted into box prompts via Grounding DINO SwinT [liu2024grounding]. For fair comparisons, all baseline-related experiments incorporate intrinsic prompts and use aligned prompt inputs.
@@
ImVoxelNet [rukhovich2022imvoxelnet] 23.5 23.4 21.5 - - - - - - 9.4
SMOKE [liu2020smoke] 25.9 20.4 20.0 - - - - - - 10.4
OV-Uni3DETR [wang2023uni3detr] 35.1 33.0 31.6 - - - - - - -
Cube R-CNN [brazil2023omni3d] 36.0 32.7 31.9 32.50 30.06 15.33 41.73 50.84 7.48 23.26
@@
$\text{DetAny3D (ours)}_{\textit{w/}~\text{Cube RCNN}}$ 35.8 33.9 32.2 31.61 30.97 18.96 46.13 54.42 7.17 24.92
@@
$\text{DetAny3D (ours)}_{\textit{w/}~\text{Ground Truth}}$ 38.0 36.7 35.9 38.68 37.55 46.14 50.62 56.82 15.98 34.38
@@ 4.2 Main Results
Zero-shot Category Performance. In this experiment, we use two sources for the prompt input: text prompt processed by Grounding DINO and box prompt from ground-truth 2D bounding box. We evaluate our model on KITTI, SUNRGBD, and ARKitScenes datasets with the same zero-shot categories as OVMono3D [yao2024open]. As shown in Table 1 (left), our DetAny3D demonstrates superior zero-shot adaptation performance compared to the OVMono3D baseline. When using Grounding DINO for text prompt input, our method achieves significant improvements of 21.02 $\rm AP_{3D}$} on KITTI, 4.29 $\rm AP_{3D}$} on SUNRGBD, and 11.35 $\rm AP_{3D}$} on ARKitScenes under the target-aware metric. When using 2D ground-truth as box prompt input, DetAny3D attains 28.96 $\rm AP_{3D}$} on KITTI, 39.09 $\rm AP_{3D}$} on SUNRGBD, and 57.72 $\rm AP_{3D}$} on ARKitScenes, showing 3.4×, 2.3×, and 4.1× gains over the baseline, respectively. This substantial performance gap highlights our method’s enhanced ability to generalize to novel object categories.
Zero-shot Camera Performance. To assess robustness against novel camera parameters, we conduct cross-dataset evaluation as shown in Table 1 (right). For Cityscapes3D and Waymo, We use Cube R-CNN’s 2D detections and ground-truth as box prompt and Grounding DINO processed text prompt for comparison. For 3RScan, due to namespace inconsistency with Cube R-CNN’s predefined categories and the presence of novel classes, we only use text prompt and ground-truth box prompts, benchmarking against OVMono3D. DetAny3D exhibits strong adaptation to unseen camera configurations. When using Cube R-CNN-aligned prompts, our model achieves $\rm{AP_{3D}}$ scores of 10.33 and 15.17 on Cityscapes3D and Waymo, respectively, surpassing Cube R-CNN by +2.11 and +5.74. With text prompts, under identical settings as OVMono3D [yao2024open], our method improves $\rm{AP_{3D}}$ by +4.73 on Cityscapes3D, +5.68 on Waymo, and +1.1 on 3RScan under target-aware metrics. Both models show low scores on conventional metrics for 3RScan due to severe naming ambiguity and missing annotations. Using 2D ground-truth as box prompts, DetAny3D attains $\rm{AP_{3D}}$ of 16.88, 15.83, and 21.36 across the three datasets, outperforming OVMono3D by +6.82, +5.6, and +3.31, respectively. These results highlight the effectiveness of our architecture and its potential for real-world applications with arbitrary camera configurations.
@@
Figure 4: Qualitative Results. We present qualitative examples from open-world detection. In each pair of images, the top row is produced by OVMono3D, and the bottom row by DetAny3D. For each example, the left sub-figure overlays the projected 3D bounding boxes, while the right sub-figure shows the corresponding bird’s-eye view with $\times1m$ grids as the background.
In-domain Performance We also evaluate our model’s in-domain detection capability using two prompt sources: 2D detections from Cube R-CNN and 2D ground-truth. Besides the unified model, we also train DetAny3D on Omni3D_out for comparison. As shown in Table 2, DetAny3D achieves competitive results with Cube R-CNN when provided with aligned input. Using GT prompts, DetAny3D outperforms OVMono3D by 9.06 $\rm{AP_{3D}}$, indicating that Cube R-CNN may bottleneck performance, and stronger 2D prompts could further boost results.
@@ 4.3 Possible Applications of DetAny3D
3D Bounding Box Guided Video Generation. We feed DetAny3D outputs into Sora for zero-shot, open-world 3D box guided video generation. As shown in Figure 3, we compare: (i) image + 3D box + text, (ii) image + 2D box + text, and (iii) image + text. With 3D box constraints, Sora generates videos better aligned with intent.
@@ 4.4 Ablation Studies
Table 3: Ablation study of DetAny3D. The table shows the impact of different design choices on $\rm AP_{3D}$} performance. Each component is progressively added. To save resources, ablations are conducted on 10% of the full training dataset.
@@
As shown in Table 3, we ablate key components of DetAny3D, showing the evolution from a SAM-based baseline to DetAny3D with strong 3D generalization. The base model extends SAM with 3D box tokens and a 3D head for direct box prediction. Additional ablations, including backbone and prompt types, are in Supplementary Section 9.
@@
  • Effectiveness of ZEM. ZEM mechanism integrates geometric features through zero-initialized layers, which enables stable 2D-to-3D knowledge transfer during training across datasets with diverse camera parameters, scenes, and depth distributions.
@@ 4.5 Qualitative Results
We provide qualitative comparisons with OVMono3D. As shown in Figure 4, our model predicts more accurate intrinsics when the camera parameters are unknown and infers more consistent camera poses and 3D detections.
@@ Acknowledgements
We sincerely thank Jiazhi Yang, Tianyu Li, Haochen Tian, Jisong Cai, and Li Chen for their invaluable discussions and constructive feedback throughout this project. Their insights and expertise have contributed significantly to the success of this work. We also appreciate the continuous support and encouragement from all the members of OpenDriveLab. This work is supported by the National Key Research and Development Program of China (2024YFE0210700), the National Natural Science Foundation of China (NSFC) under Grants 62206172 and 62432008, and the Shanghai Artificial Intelligence Laboratory. It is also partially funded by Meituan Inc.
[…]
@@ 6 DA3D
DA3D is a unified 3D detection dataset, consists of 16 diverse datasets. It builds upon six datasets in Omni3D—Hypersim [roberts2021hypersim], ARKitScenes [baruch1arkitscenes], Objectron [ahmadyan2021objectron], SUNRGBD [song2015sun], KITTI [geiger2013vision], and nuScenes [caesar2020nuscenes]—while partially incorporating an additional 10 datasets to further enhance the scale, diversity, and generalization capabilities of 3D detection models. As shown in Figure 5, DA3D comprises 0.4 million frames (2.5× the scale of Omni3D), spanning 20 distinct camera configurations.
The dataset is standardized with the similar structure to Omni3D [brazil2023omni3d], including monocular RGB images, camera intrinsics, 3D bounding boxes, and depth maps. DA3D is designed to test 3D detection models across a wide variety of environments, camera configurations, and object categories, offering a more comprehensive evaluation setting.
@@ 6.1 Dataset Composition
For training and evaluation, we follow the dataset splitting strategy used in prior works [brazil2023omni3d]. Specifically:
@@ 6.3 Evaluation Setup
Zero-Shot Categories. Following prior work [yao2024open], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as unseen classes for zero-shot testing.